OpenStudy (anonymous):

If Logx (1 / 8) = - 3 / 2, then x is equal to A. - 4 B. 4 C. 1 / 4 D. 10

OpenStudy (anonymous):

\[x^{-\frac{3}{2}}=\frac{1}{8}\] \[x^{\frac{3}{2}}=8\] \[x^3=64\] \[x=\sqrt[3]{64}=4\]

OpenStudy (anonymous):

initial step is important one \[\log_b(x)=y\iff b^y=x\]

OpenStudy (anonymous):

thanks!!! that helps a lot!!!!!:D